# DDPG-E2E: A Novel Policy Gradient Approach for End-to-End Communication Systems

### v1

@@ IV-A Deep Reinforcement Learning Basics
where $G_{t}=\sum^{\infty}_{t=1}\gamma^{t}r_{t}$ denotes the cumulative discounted reward for one trajectory, and $\gamma\in(0,1]$ represents the discount factor which determines the importance of future rewards relative to immediate rewards. In Q-learning, the update rule of the Q-value used for learning the optimal policy can be expressed as:
@@
$\begin{split}Q_{\pi}(\bm{s},\pi^{\prime}(\bm{s}))=\frac{\epsilon}{|A|}\sum_{\bm{a}\in\mathbb{A}}Q_{\pi}(\bm{s},\bm{a})+(1-\epsilon)Q_{\pi}(\bm{s},\bm{a}),\end{split}$ (4)
where $\epsilon$ is the probability to make exploration at every time step. In practice, the probability of exploration $\epsilon$ will be decreased gradually according to the convergence situation.
To further enhance the convergence of Q-learning, the concept of deep Q-learning has been introduced, replacing the traditional Q-table with a DNN, known as deep Q-network (DQN) [20]. DQN harnesses the computational power of DNN to address regression problem by taking the state-action pair as input and approximating the corresponding Q-value through predictive modelling. The primary advantage of deep Q-learning over conventional Q-learning lies in its efficiency, particularly when dealing with complex tasks in which Q-learning requires very long time to learn or even cannot obtain the optimal solution. Furthermore, deep Q-learning introduces an innovative technique known as experience replay [21]. This method involves storing accumulated information in a buffer, enabling the agent to sample from the past experience to optimize the policy network. The experience replay effectively utilizes the past cumulative experience and mitigates the issues of data correlation in the observation sequences by random sampling. However, deep Q-learning encounters several limitations. The most critical issue of deep Q-learning is that it cannot directly operate in continuous actions space, whereas many real-world applications precisely require continuous actions space. For instance, practical communication systems are often involved with continuous, time-varying encoded signal and can be difficult to measure by discrete states. Another critical challenge is the high update frequency of target network, leading to the behaviour network continuously “chasing” a moving target [22]. To overcome these challenges, this article introduces a DDPG-based approach for E2E communication systems. This proposed scheme not only handle the problem of continuous action space, but also enables joint optimizations on both the transmitter and the receiver without relying on prior channel information.
@@ IV-B End-to-End Communication Systems using DDPG
DDPG, or Deep Deterministic Policy Gradient, represents an off-policy actor-critic algorithm [23] that effectively combines the advantages of deep Q-learning and policy gradient [24]. The proposed DDPG-based system encompasses several significant features: (i) it employs a deterministic policy for network optimizations, allowing the agent to directly operate on the continuous action space, (ii) the system adopts actor-critic approach, where the actor network is responsible for generating the deterministic action based on the current state, and the critic network provides a score that evaluates the performance of the current action, and (iii) to enable training stability, the DDPG approach utilizes two target networks - one for the actor and another for the critic. These target networks serve as time-delayed copies of their original networks that slowly track the behaviour of the behaviour networks. This off-policy design significantly contributes to the overall stability of the training process.
@@
The schematic overview of the proposed DDPG-based E2E communication system is visualized in Fig. 2. The observation state corresponds to the input message, i.e., $\bm{s}_{t}=\bm{m}$. The action corresponds to the encoded signal, i.e., $\bm{a}_{t}=\bm{x}$. The received signal is represented as $\bm{y}$, while $\hat{\bm{m}}$ signifies the estimated output message. Additionally, we have $\bm{s}_{\mathrm{b}}$ and $\bm{a}_{\mathrm{b}}$, which denote the randomly sampled state and action batches from the experience buffer used for training the transmitter and the receiver. At the core of the system, the environment comprises two key elements, including the channel model and the receiver model. The DDPG agent encompasses the actor and the critic networks. The actor network, serving as the transmitter model, maps the input message $\bm{m}$ to the encoded signal $\bm{x}$ and transmits it through the unknown channel. The channel model introduces distortions to the encoded signal $\bm{x}$ before sending it to the receiver model. The receiver model then decodes and recovers the distorted signal to produce the estimated message $\hat{\bm{m}}$. The binary cross-entropy loss [25] is employed to measure the distance between the original information $\bm{m}$ and the estimated message $\hat{\bm{m}}$. The immediate reward after performing an action is the opposite of the computed loss since the loss negatively reflects the model performance, which can be expressed as:
$r_{t}=$ $-\sum\limits_{i=1}^{K}((\bm{s}_{t})_{i}[\log(f_{\bm{\theta}_{\mathrm{R}}}(\mu(\bm{s}_{t}\mid\bm{\theta}_{\mathrm{T}})))]_{i}$ (5)
$+(1-(\bm{s}_{t})_{i})[\log(1-f_{\bm{\theta}_{\mathrm{R}}}(\mu(\bm{s}_{t}\mid\bm{\theta}_{\mathrm{T}})))]_{i}),$
where $\bm{\theta}_{\mathrm{R}}$ denotes the weights of the receiver network, $\bm{\theta}_{\mathrm{T}}$ denotes the weights of the transmitter network, and $\mu(\bm{s}_{t}|\bm{\theta}_{\mathrm{T}})$ denotes the action taken by policy $\mu$ at state $\bm{s}_{t}$ based on the transmitter weights $\bm{\theta}_{\mathrm{T}}$. $f_{\bm{\theta}_{\mathrm{R}}}$ represents the process at the receiver model mapping the distorted signal to the estimated message, and $K$ denotes the length of the input message. The reward $r_{t}$ is the feedback signal from the receiver for updating both the actor and the critic networks at the agent.
@@ IV-C Network Architectures of Transmitter, Receiver and Critic
Fig. 3 provides an illustration of the network architectures of the transmitter, the receiver, and the critic models. In our proposed solution, CNNs are employed to facilitate the curse of the dimensionality. The detailed parameter configurations for the network architectures of the transmitter and the receiver are presented in Table I, with $K$ denoting the block length. At the transmitter, the original message $\bm{m}$ is a binary number vector of size $K$, i.e., $\bm{m}\in\{0,1\}^{K}$. This message $\bm{m}$ serves as the input to the transmitter. More specifically, the transmitter architecture depicted in Fig. 3(a) comprises four one-Dimentional convolutional (Conv1D) layers. Subsequently, an L2 normalization layer [26] is applied to normalize the total power of the encoded signal, ensuring $||\bm{x}||^{2}\leqslant n$. The initial convolutional layer employs 256 filters, followed by a batch normalization layer and the Exponential Linear Unit (ELU) activation function [27]. The number of filters progressively decreases to 2 for reshaping the dimension of the encoded signal. The resulting encoded signal would be a two-dimensional vector with a length of $K$, representing the complex-valued symbols.
At the receiver, the estimated channel response $\hat{\bm{h}}$ and the received signal $\bm{y}$ are combined and served as the inputs of the receiver network. The receiver network is composed of three Conv1D layers, followed by an additional Conv1D layer with Sigmoid activation function [28]. The latter layer is responsible for generating the estimates of the probabilities associated with all possible output messages $\bm{p}$. This output message vector $\bm{p}$ will be rounded to a binary number vector, i.e., the estimated message $\hat{\bm{m}}$. The block rint shown in Fig. 3(b) denotes that probabilities of the output vector are rounded to integers. Fig. 3(c) shows the network architecture of the critic model. The critic network takes both the message $\bm{m}$ and the encoded signal $\bm{x}$ as the input, and generates $Q(\bm{m},\bm{x})$ as the critic value, where each input follows with a flatten layer and a dense layer. The aligned inputs are then concatenated to a one-dimension vector with a length of 128. The concatenated input is then passed through three sequential layers, each consisting of a fully connected (Dense) layer followed by Batch Normalization [29] and Rectified Linear Unit (ReLU) activation function [30]. The first two Dense layers consist of 256 neurons each, while the final Dense layer has a single neuron output for generating the critic value. The final Dense layer is followed by Batch Normalization and a linear activation function.
@@
Conv1D+BN+ELU 5 $K\times 256$
Conv1D+BN+ELU 3 $K\times 128$
Conv1D+BN+ELU 3 $K\times 64$
@@
Conv1D+BN+ELU 5 $K\times 256$
Conv1D+BN+ELU 5 $K\times 128$
Conv1D+BN+ELU 5 $K\times 64$
@@
Dense1+ReLU N/A 64
Dense2+ReLU N/A 64
Concatenate N/A 128
@@ V-A Training Process Overview
Observe the new state $\bm{s}_{t+1}$, reward $r_{t}$ and done information
@@
13:
@@
$L=\frac{1}{B}\sum_{i}(\bm{y}_{i}-Q(\bm{s}_{i},\bm{a}_{i}|\bm{\theta}^{Q}))^{2}$
16:
@@
17:         Update Target networks for both actor and critic:
@@
18:
Sample a random batch of $D$ transitions $(\bm{a}_{t},\bm{s}_{t})$
19:
Train Receiver model with the sampled batch
20:         if done==True then
21:              Break
22:         end if
23:     end for
24:     if stop criterion met then
25:         Break Main Training Loop
26:     end if
27:end for
@@
The diagram overview of the training process is depicted in Fig. 2 and the pseudocode of the whole training process is presented in Algorithm 1. At the initial phase, both the actor and the critic networks are initialized with random parameters $\bm{\theta}^{\mu}$ and $\bm{\theta}^{Q}$, respectively. Two target networks, denoted as $\mu^{\prime}$ and $Q^{\prime}$ are initialized with identical parameters for the initial equilibrium. A replay buffer with a capacity of $C$ is initialized to store the cumulative information obtained from the environment. During the training phase, a random observation state is generated at the beginning of each episode, which can be represented as $\bm{s}_{t}=\bm{m}$. At every time step within each episode, the actor network maps the current observation state $\bm{s}_{t}$ to the action, i.e., encoded signal $\bm{a}_{t}=\bm{x}$. The chosen action is then conveyed to the environment, where the observation state is considered as the label of the estimated output for calculating the loss value. Subsequently, the environment will then return the next observation state $\bm{s}_{t+1}$, the calculated reward $r_{t}$ and the $done$ information as feedback to the agent, where the $done$ information is a Boolean variable that determines whether to terminate the training of current episode. A collection of state, action, reward, and next state will be stored in the experience buffer of the agent. Once the data is collected, a mini-batch of experiences is randomly sampled from the replay buffer. For each state-action pair in the mini-batch, the target Q-value is computed using the following equation:
$\bm{y}_{i}$ $=r_{i}+\gamma{Q(\bm{s}_{i+1},\bm{a}_{i+1})},$ (6)
$=r_{i}+\gamma{Q(\bm{s}_{i+1},\mu(\bm{s}_{i+1}|\bm{\theta}^{\mu^{\prime}})|\bm{\theta}^{Q^{\prime}})}.$
@@
The critic network is then trained to minimize the mean square error between the predicted Q-value and the target Q-value using the mini-batch of experiences. The loss function between the estimated Q-value and the target Q-value can be expressed as:
$L=\frac{1}{B}\sum_{i}(\bm{y}_{i}-Q(\bm{s}_{i},\bm{a}_{i}|\bm{\theta}^{Q}))^{2},$ (7)
@@
where $Q(\bm{s}_{i},\bm{a}_{i}|\bm{\theta}^{Q})$ denotes the predicted Q-value generated by the behaviour networks. In the process of optimizing the actor and the critic networks, the objective is to iteratively update these networks to maximize the expected return, i.e., the the expected value of the Q-function. The expected value of the Q-function can be mathematically represented as:
@@
where $\bm{s}_{i}$ denotes the current state of the environment, and $\bm{a}_{i}$ denotes the current action selected by the policy function, i.e., $\bm{a}_{i}=\mu(\bm{s}_{i})$. In the optimization process of the actor network, a crucial step is to calculate the gradient of the expected Q-value concerning the policy parameters $\bm{\theta}_{\mathrm{\mu}}$. This enables the actor network to learn to improve its policy, leading to more favorable actions in different states. The gradient of the expected Q-value with respect to the policy parameter is mathematically formulated as:
@@
where $\bm{\theta}^{\prime}$ represents the parameters of target networks, $\bm{\theta}$ denotes the parameters of behaviour networks, and $\tau$ is a hyper-parameter referred to as “soft update rate”. This parameter controls the gradual adjustments of target network parameters to align with those of the behaviour networks, preventing abrupt policy changes during training. $\tau$ is usually defined as a positive value that is much smaller than 1 for slow update, i.e., $\tau\ll 1$. Furthermore, the receiver model undergoes supervised learning with the state-action pair batches randomly sampled from the buffer as the output data and input data, respectively. This training approach allows our proposed solution to optimize both the transmitter and the receiver without requiring prior knowledge of the channel model.

### v2

@@ IV-A Deep Reinforcement Learning Basics
where $G_{t}=\sum^{\infty}_{t=1}\gamma^{t}r_{t}$ denotes the cumulative discounted reward for one trajectory, and $\gamma\in(0,1]$ represents the discount factor which determines the importance of future rewards relative to immediate rewards. In Q-learning, the Q-value update rule for learning the optimal policy is given by:
@@
$\begin{split}Q_{\pi}(\bm{s},\pi^{\prime}(\bm{s}))=\frac{\epsilon}{|A|}\sum_{\bm{a}\in\mathbb{A}}Q_{\pi}(\bm{s},\bm{a})+(1-\epsilon)\max_{\bm{a}\in\mathbb{A}}Q_{\pi}(\bm{s},\bm{a}),\end{split}$ (4)
where $\epsilon$ is the probability of making an exploratory action at every time step. In practice, the exploration probability $\epsilon$ is typically decreased gradually as the algorithm converges, allowing the policy to exploit the learned information more often. $\pi^{\prime}$ refers to the target policy, which selects actions based on the balance between exploration and exploitation, while $\pi$ refers to the current policy being evaluated.
To further enhance the convergence of Q-learning, the concept of deep Q-learning has been introduced, replacing the traditional Q-table with a NN, known as deep Q-network (DQN) [24]. DQN harnesses the computational power of NN to address regression problems by taking the state-action pair as input and approximating the corresponding Q-value through predictive modelling. The primary advantage of deep Q-learning over conventional Q-learning lies in its efficiency, particularly when dealing with complex tasks in which Q-learning requires a very long time to learn or even cannot obtain the optimal solution. Furthermore, deep Q-learning introduces an innovative technique known as experience replay [25]. This method involves storing accumulated information in a buffer, enabling the agent to sample from past experiences to optimize the policy network. The experience replay effectively utilizes the past cumulative experience and mitigates the issues of data correlation in the observation sequences by random sampling. However, deep Q-learning encounters several limitations. The most critical issue of deep Q-learning is that it cannot directly operate in continuous actions space, whereas many real-world applications precisely require continuous actions space. For instance, practical communication systems are often involved with continuous, time-varying encoded signals and can be difficult to measure by discrete states. Another challenge is the frequent update of the target network, causing the behaviour network continuously ”chasing” a moving target [26]. To address this, we propose a DDPG-based approach for E2E communication systems. This method effectively manages continuous action spaces and allows joint optimization for the transceivers without requiring prior channel information.
@@ IV-B End-to-End Communication Systems using DDPG
DDPG, or Deep Deterministic Policy Gradient, represents an off-policy actor-critic algorithm [27] that effectively combines the advantages of deep Q-learning and policy gradient [28]. The proposed DDPG-based system encompasses several key features: (i) it employs a deterministic policy for network optimizations, allowing the agent to directly operate on the continuous action space, (ii) it follows an actor-critic architecture, where the actor network generates the deterministic action based on the current state, and the critic network evaluates the performance of the current action, and (iii) to ensure stable training, the DDPG approach uses two target networks - one for the actor and another for the critic. These target networks serve as time-delayed copies of their original networks that slowly track the behaviour of the behaviour networks. This off-policy design significantly contributes to the overall stability of the training process.
@@
The schematic overview of the proposed DDPG-based E2E communication system is visualized in Fig. 2. The observation state corresponds to the input message at time step $t$, i.e., $\bm{s}_{t}=\bm{m}_{t}$. The action represents the encoded signal, i.e., $\bm{a}_{t}=\bm{x}_{t}$. The received signal is denoted by $\bm{y}_{t}$, while $\hat{\bm{m}}$ signifies the estimated output message. Additionally, $\bm{s}_{\mathrm{b}}$ and $\bm{a}_{\mathrm{b}}$ represent the randomly sampled state and action batches from the experience buffer used for training the transmitter and receiver. At the core of the system, the environment comprises two key elements, including the channel model and the receiver model. The DDPG agent encompasses the actor and the critic networks. The actor network, serving as the transmitter model, maps the input message $\bm{m}$ to the encoded signal $\bm{x}$ and transmits it through the unknown channel. The channel model introduces distortions to the encoded signal $\bm{x}$ before sending it to the receiver model. The receiver model then decodes and recovers the distorted signal to produce the estimated message $\hat{\bm{m}}$. The binary cross-entropy loss [29] is employed to measure the distance between the original information $\bm{m}$ and the estimated message $\hat{\bm{m}}$. The immediate reward after performing an action is the opposite of the computed loss since the loss negatively reflects the model performance, which can be expressed as:
$r_{t}=$ $-\sum\limits_{k=1}^{K}\left[(\bm{s}_{t})_{k}\log\left(f_{\bm{\theta}_{\mathrm{R}}}\left(\mu(\bm{s}_{t}\mid\bm{\theta}_{\mathrm{T}})\right)\right)_{k}\right.$ (5)
$\left.+\left(1-(\bm{s}_{t})_{k}\right)\log\left(1-f_{\bm{\theta}_{\mathrm{R}}}\left(\mu(\bm{s}_{t}\mid\bm{\theta}_{\mathrm{T}})\right)\right)_{k}\right],$
where $\bm{\theta}_{\mathrm{R}}$ denotes the weights of the receiver network, $\bm{\theta}_{\mathrm{T}}$ denotes the weights of the transmitter network, and $\mu(\bm{s}_{t}|\bm{\theta}_{\mathrm{T}})$ denotes the action taken by policy $\mu$ at state $\bm{s}_{t}$ based on the transmitter weights $\bm{\theta}_{\mathrm{T}}$. $f_{\bm{\theta}_{\mathrm{R}}}$ represents the process at the receiver model mapping the distorted signal to the estimated message, and $K$ denotes the length of the input message. The reward $r_{t}$ is the feedback signal from the receiver for updating both the actor and the critic networks at the agent. $k$ denotes the number of indices in the binary message. Due to the dynamic nature of wireless communications, the transmitted bits are typically independent and uncorrelated, resembling a multi-armed bandit problem. To align with the randomness between the signals, we assume that the adjacent states are uncorrelated and independent of each other. A similar approach has been implemented in [17], where the author also assumes the uncorrelated input batches. DRL enables the system to optimize the transceivers by leveraging an exploration mechanism, thereby facilitating adaptation from known to unknown channel conditions. Since the next state $\bm{s}_{t+1}$ is randomly generated and is independent of the current state $\bm{s}_{t}$, the return of the Bellman function should focus more on the immediate reward rather than the future value. For that, the discount factor $\gamma$ is set at 0.01 to eliminate the correlated effect between the adjacent states [30].
@@ IV-C Network Architectures of Transmitter, Receiver and Critic
Fig. 3 provides an illustration of the network architectures of the transmitter, the receiver, and the critic models. In our proposed solution, CNNs are employed to facilitate the curse of dimensionality. The detailed parameter configurations for the network architectures of the transmitter and the receiver are presented in Table I, with $K$ denoting the block length. At the transmitter, the original message $\bm{m}$ is a binary number vector of size $K$, i.e., $\bm{m}\in\{0,1\}^{K}$. This message $\bm{m}$ serves as the input to the transmitter. More specifically, the transmitter architecture depicted in Fig. 3(a) comprises four one-Dimentional convolutional (Conv1D) layers using circular padding techniques [20]. Subsequently, an L2 normalization layer [31] is applied to normalize the total power of the encoded signal, ensuring $||\bm{x}||^{2}\leqslant n$. The initial convolutional layer employs 256 filters, followed by a batch normalization (BN) layer [32] and the Mish activation function [33]. The number of filters progressively decreases to 2 for reshaping the dimension of the encoded signal. The resulting encoded signal would be a two-dimensional vector with a length of $K$, representing the complex-valued symbols.
At the receiver, the estimated channel response $\hat{\bm{h}}$ and the received signal $\bm{y}$ are combined and served as the inputs of the receiver network. The receiver network is composed of three Conv1D layers, followed by an additional Conv1D layer with the Sigmoid activation function [34]. The latter layer is responsible for generating the estimates of the probabilities associated with all possible output messages $\bm{p}$. This output message vector $\bm{p}$ will be rounded to a binary number vector, i.e., the estimated message $\hat{\bm{m}}$. The block rint shown in Fig. 3(b) denotes that probabilities of the output vector are rounded to integers. Fig. 3(c) shows the network architecture of the critic model. The critic network takes both the message $\bm{m}$ and the encoded signal $\bm{x}$ as the input and generates $Q(\bm{m},\bm{x})$ as the critic value, where each input follows with a Flatten layer and a dense layer. The aligned inputs are then concatenated to a one-dimension vector with a length of 256. The concatenated input is then passed through three sequential layers, each consisting of a fully connected (Dense) layer followed by BN and Rectified Linear Unit (ReLU) activation function [35]. The first two Dense layers consist of 256 neurons each, while the final Dense layer has a single neuron output for generating the critic value. The final Dense layer is followed by BN and a linear activation function.
@@
Conv1D+BN+Mish 5 $K\times 256$
Conv1D+BN+Mish 3 $K\times 128$
Conv1D+BN+Mish 3 $K\times 64$
@@
Conv1D+BN+Mish 5 $K\times 256$
Conv1D+BN+Mish 5 $K\times 128$
Conv1D+BN+Mish 5 $K\times 64$
Conv1D+BN+Mish 5 $K\times 32$
@@
Dense1+ReLU N/A 128
Dense2+ReLU N/A 128
Concatenate N/A 256
@@ V-A Training Process Overview
Observe the new state $\bm{s}_{t+1}$, reward $r_{t}$
@@
13:         TrainAgent
14:         TrainRx
15:     end for
16:end for
Algorithm 2 Training of the Agent (Actor, Critic, Target Actor and Target Critic)
1:function TrainAgent()
2:     Input:
@@
$\mathcal{L}_{\bm{\theta}^{\mathrm{\mu}}}=\frac{1}{B}\sum_{i}(\bm{y}_{i}^{\text{target}}-Q(\bm{s}_{i},\bm{a}_{i}|\bm{\theta}^{Q}))^{2}$
5:
@@
6:     Update Target networks for both actor and critic:
@@
7:end function
Algorithm 3 Training of the Receiver
1:function TrainRx
2:     Input: Sample state-action batches $(\bm{s}_{b},\bm{a}_{b})$ with a batch size of $B$
3:     Perform supervised learning with the sampled state batch as the input data, and the sampled action batch as the label
4:
$f_{\bm{\theta}_{\mathrm{R}}}({\bm{a}_{b}})={\hat{\bm{s}}_{b}}$
5:     Calculate the binary cross-entropy loss for the receiver
$\mathcal{L}_{\bm{\theta}_{\mathrm{R}}}=-\frac{1}{K}\sum_{k=1}^{K}{(\bm{s}_{b})}_{k}\log({(\hat{\bm{s}}_{b})}_{k})+(1-{(\bm{s}_{b})}_{k})\log(1-{(\hat{\bm{s}}_{b})}_{k})$
6:     Update the model parameters using the computed loss
7:end function
@@
The diagram overview of the training process is depicted in Fig. 2 and the pseudocode of the whole training process is presented in Algorithm 1. At the initial phase, both the actor and the critic networks are initialized with random parameters $\bm{\theta}^{\mu}$ and $\bm{\theta}^{Q}$, respectively. Two target networks, denoted as $\mu^{\prime}$ and $Q^{\prime}$ are initialized with identical parameters for the initial equilibrium. A replay buffer with a capacity of $C$ is initialized to store the cumulative information obtained from the environment. During the training phase, a random observation state is generated at the beginning of each episode, which can be represented as $\bm{s}_{t}=\bm{m}$. At every time step within each episode, the actor network maps the current observation state $\bm{s}_{t}$ to the action, i.e., encoded signal $\bm{a}_{t}=\bm{x}$. The chosen action is then conveyed to the environment, where the observation state is considered as the label of the estimated output for calculating the loss value. Subsequently, the environment will then return the next observation state $\bm{s}_{t+1}$, the calculated reward $r_{t}$ and the $done$ information as feedback to the agent, where the $done$ information is a Boolean variable that determines whether to terminate the training of current episode. A collection of state, action, reward, and next state will be stored in the experience buffer of the agent.Once the data is collected, it starts to train the agent as depicted in Algorithm 2. A mini-batch of experiences is randomly sampled from the replay buffer. For each state-action pair in the mini-batch, the target Q-value $\bm{y}_{i}^{\text{target}}$ is computed using the following equation:
$\bm{y}_{i}^{\text{target}}$ $=r_{i}+\gamma Q(\bm{s}_{i+1},\bm{a}_{i+1})$ (6)
$=r_{i}+\gamma Q(\bm{s}_{i+1},\mu(\bm{s}_{i+1}\mid\bm{\theta}^{\mu^{\prime}})\mid\bm{\theta}^{Q^{\prime}}).$
@@
The critic network is then trained to minimize the mean square error between the predicted and target Q-values using the mini-batch of experiences. The loss function between the estimated and target Q-values can be expressed as:
$\mathcal{L}_{\bm{\theta}^{\mathrm{\mu}}}=\frac{1}{B}\sum_{i}(\bm{y}_{i}^{\text{target}}-Q(\bm{s}_{i},\bm{a}_{i}|\bm{\theta}^{Q}))^{2},$ (7)
@@
where $Q(\bm{s}_{i},\bm{a}_{i}|\bm{\theta}^{Q})$ denotes the predicted Q-value generated by the behaviour networks. In the process of optimizing the actor and the critic networks, the objective is to iteratively update these networks to maximize the expected return, i.e., the expected value of the Q-function. The expected value of the Q-function can be mathematically represented as:
@@
where $\bm{s}_{i}$ denotes the current state of the environment, and $\bm{a}_{i}$ denotes the current action selected by the policy function, i.e., $\bm{a}_{i}=\mu(\bm{s}_{i})$. In the optimization process of the actor network, a crucial step is to calculate the gradient of the expected Q-value concerning the policy parameters $\bm{\theta}_{\mathrm{\mu}}$. This enables the actor network to learn more favorable actions in different states. The gradient of the expected Q-value with respect to the policy parameter is mathematically formulated as:
@@
where $\bm{\theta}^{\prime}$ represents the parameters of target networks, $\bm{\theta}$ denotes the parameters of behaviour networks, and $\tau$ is a hyper-parameter referred to as “soft update rate”. This parameter controls the gradual adjustments of target network parameters to align with those of the behaviour networks, preventing abrupt policy changes during training. $\tau$ is usually defined as a positive value that is much smaller than 1 for slow update, i.e., $\tau\ll 1$. Furthermore, the receiver is trained using supervised learning. In this process, batches of state-action pairs are randomly sampled from the buffer. The state-action pairs are used as input, with their corresponding outputs serving as labels. The receiver network is optimized by minimizing the binary cross-entropy loss. The training approach allows our proposed solution to optimize the transceivers without requiring prior knowledge of the channel model.